# Extremal Steklov-Neumann eigenvalues

### v1

@@ Abstract.
Let $\Omega$ be a bounded open planar domain with smooth connected boundary, $\Gamma$, that has been partitioned into two disjoint components, $\Gamma=\Gamma_{S}\sqcup\Gamma_{N}$. We consider the Steklov-Neumann eigenproblem on $\Omega$, where a harmonic function is sought that satisfies the Steklov boundary condition on $\Gamma_{S}$ and the Neumann boundary condition on $\Gamma_{N}$. We pose the extremal eigenvalue problems (EEPs) of minimizing/maximizing the $k$-th non-trivial Steklov-Neumann eigenvalue among boundary partitions of prescribed measure. We formulate a relaxation of these EEPs in terms of weighted Steklov eigenvalues where an $L^{\infty}(\Gamma)$ density replaces the boundary partition. For these relaxed EEPs, we establish existence, prove optimality conditions, show that the maximization problem is convex for $k=1$ and non-convex for $k\geq 2$, and establish symmetry properties for the maximizing densities for $k=1$. We also prove a homogenization result that allows us to use solutions to the relaxed EEPs to infer properties of solutions to the original EEPs. For a disk, we provide numerical and asymptotic evidence that the minimizing arrangement of $\Gamma_{S}\sqcup\Gamma_{N}$ for the $k$-th eigenvalue consists of $k+1$ connected components that are symmetrically arranged on the boundary. For a disk, we prove that for $k=1$, the constant density is a maximizer for the relaxed problem; we also provide numerical and asymptotic evidence that for $k\geq 2$, the maximizing density for the relaxed problem is a non-trivial function; a sequence of rapidly oscillating Steklov/Neumann boundary conditions approach the supremum value.
@@ 1. Introduction
Fix a bounded open domain $\Omega\subset\mathbb{R}^{2}$ with smooth connected boundary $\partial\Omega=\Gamma$. For a disjoint partition of the boundary, $\Gamma=\Gamma_{S}\sqcup\Gamma_{N}$, we consider the Steklov-Neumann (S-N) eigenproblem:
$\Delta u$ $=0$  $\qquad\text{ in $\Omega$},$ (1a)
$\partial_{n}u$ $=\sigma u$  $\qquad\text{ on $\Gamma_{S}$},$ (1b)
$\partial_{n}u$ $=0$  $\qquad\text{ on $\Gamma_{N}$}.$ (1c)
@@
$\sigma_{k,\alpha}^{\bigtriangledown}=\inf_{\begin{subarray}{c}\Gamma_{S}\subset\Gamma\\ \left\lvert\Gamma_{S}\right\rvert=\alpha\left\lvert\Gamma\right\rvert\end{subarray}}\ \sigma_{k}(\Gamma_{S}).$ (2)
@@
$\sigma_{k,\alpha}^{\triangle}=\sup_{\begin{subarray}{c}\Gamma_{S}\subset\Gamma\\ \left\lvert\Gamma_{S}\right\rvert=\alpha\left\lvert\Gamma\right\rvert\end{subarray}}\ \sigma_{k}(\Gamma_{S}).$ (3)
@@
The optimization problems (2) and (3) were posed by Nilima Nigam in the AIM Problem List (AimPL) on Steklov eigenproblems [22, Problem 4.1]. Steklov eigenvalues appear in many applications [21, 13], so the extremal eigenvalue problems (2) and (3) are of fundamental interest. We offer a physical interpretation in terms of the analogous extremal Laplacian eigenvalue problems with mixed Dirichlet/Neumann boundary conditions (B.C.) studied by S. J. Cox and P. X. Uhlig [10]. These problems, for $k=1$, have the interpretation that if we are allowed to fasten a prescribed fraction of the boundary of a drum (corresponding to Dirichlet B.C.), which portion should we fasten to produce the lowest/highest bass note (first Laplacian eigenvalue)? The S-N eigenvalue problem (1) is the limit of this eigenvalue problem where the mass of the drum is concentrated at $\Gamma_{S}$ [23]. Thus, the optimization problem (2) for $k=1$ has the following physical interpretation: Suppose we are allowed to fasten a fixed fraction of the boundary of a drum (corresponding to Steklov B.C.), where the mass of the drum will also be concentrated. Which portion should we fasten to produce the lowest bass note?
@@
The S-N eigenvalue problem has attracted the interest of researchers in recent years. In [6], Bañuelos, Kulczycki, Polterovich, and Siudeja established inequalities between mixed S-N and mixed Steklov-Dirichlet eigenvalues for certain domains satisfying the weak John’s condition. These inequalities are used to obtain geometric information about the nodal sets of S-N eigenfunctions. In [4], Ammari, Imeri, and Nigam analyzed how small changes in the S-N partition move the spectrum and used the resulting shape-derivative formulas to design a boundary-integral method that tunes mixed S-N conditions to hit target resonances. They also treated multiple eigenvalues and validated the asymptotics numerically on canonical shapes. In [5], Arias-Marco, Dryden, Gordon, Hassannezhad, Ray, and Stanhope proved a sharp isoperimetric inequality for mixed Steklov eigenvalues on surfaces by relating S-N and Steklov–Dirichlet spectra, developed full asymptotics for mixed problems under mild interface regularity, and studied symmetry-constrained maximization, showing that the upper bounds converge to the Hersch-Payne-Schiffer asymptotics as $k\to\infty$. In [14], Grebenkov gave a “small-window” asymptotic analysis for the mixed S-N problem (framed via an integral-operator reformulation). The asymptotics are checked numerically for an arc on a disk and a spherical cap on a ball, and are linked to diffusion-controlled reactions (mean first-reaction times and reactivity-dependent corrections). In [8], Basak and Verma studied Steklov and mixed S-N spectra on domains with a spherical hole. They obtained an explicit expression for the second nonzero Steklov eigenvalue on an annulus and derived sharp upper bounds for the first $n$ nonzero Steklov eigenvalues on 4–fold symmetric domains in $\mathbb{R}^{n}$, with analogous bounds for the mixed problem; examples show the symmetry hypothesis is essential.
@@
$\Delta u$ $=0$  $\qquad\text{ in $\Omega$},$ (4a)
$\partial_{n}u$ $=\lambda\rho u$  $\qquad\text{ on $\Gamma$}.$ (4b)
@@
$\lambda_{k}=\min_{E_{k}\subset H^{1}(\Omega)}\ \max_{u\in E_{k}\setminus\{0\}}\ \frac{\int_{\Omega}\left\lvert\nabla u\right\rvert^{2}dx}{\int_{\Gamma}\rho u^{2}\,ds}.$ (5)
This weighted Steklov eigenproblem is of interest in its own right, as a conformal mapping applied to the Steklov eigenvalue problem will introduce a boundary weight [31, 2, 24].
@@
$\mathcal{A}_{\alpha}\coloneqq\left\{\rho\in L^{\infty}(\Gamma)\colon 0\leq\rho(s)\leq 1\textrm{ for a.e. }s\in\Gamma,\,\int_{\Gamma}\rho(s)\,ds=\alpha\left\lvert\Gamma\right\rvert\right\}$. (6)
@@
$\lambda_{k,\alpha}^{\bigtriangledown}=\inf_{\rho\in\mathcal{A}_{\alpha}}\ \lambda_{k}(\rho),$ (7)
@@
$\lambda_{k,\alpha}^{\triangle}=\sup_{\rho\in\mathcal{A}_{\alpha}}\ \lambda_{k}(\rho).$ (8)
Clearly, we have that $\lambda_{k,\alpha}^{\bigtriangledown}\leq\sigma_{k,\alpha}^{\bigtriangledown}\leq\sigma_{k,\alpha}^{\triangle}\leq\lambda_{k,\alpha}^{\triangle}$. Note that these are all finite, since by the Hersch-Payne-Schiffer inequality [16, 12]
$\lambda_{k}(\rho)\leq\frac{2\pi}{\alpha\left\lvert\Gamma\right\rvert}k,\qquad\forall\rho\in\mathcal{A}_{\alpha}.$ (9)
@@
(Existence) For every $k\in\mathbb{N}$ and $\alpha\in(0,1]$, there exists $\rho_{k}^{\bigtriangledown}\in\mathcal{A}_{\alpha}$ that attains the infimum in (7) and there exists $\rho_{k}^{\triangle}\in\mathcal{A}_{\alpha}$ that attains the supremum in (8). For $k=1$ and $\alpha\in(0,1]$, there exists a “bang-bang” density $\rho_{1}^{\bigtriangledown}\in\mathcal{A}_{\alpha}$ such that $\lambda_{1}(\rho_{1}^{\bigtriangledown})=\lambda_{1,\alpha}^{\bigtriangledown}=\sigma_{1,\alpha}^{\bigtriangledown}$.
Theorem 1.1 is proven in Section 2. The first statement follows from the continuity of $\rho\mapsto\lambda_{k}(\rho)$ on $\mathcal{A}_{\alpha}$ and the compactness of $\mathcal{A}_{\alpha}$ with respect to the weak topology on $L^{\infty}(\Gamma)$. Here, we say that $\rho\in\mathcal{A}_{\alpha}$ is a “bang-bang” density if $\rho=\mathds{1}_{S}$ for a measurable set $S\subseteq\Gamma$. To prove a minimizing density is bang-bang for $k=1$, we establish the convexity of the map $\rho\mapsto\lambda_{1}^{-1}(\rho)$; see Lemma 2.2(c).
@@
For $k\geq 2$, $\rho\mapsto\lambda_{k}^{-1}(\rho)$ is not convex; see Remark 3.1. This seems to be misunderstood in the literature for eigenvalues of other elliptic operators. In Appendix A, we give a simple example of a Sturm-Liouville operator that has non-convex higher eigenvalues.
@@
(Homogenization) Suppose $\rho\in L^{\infty}(\Gamma)$ and $\rho_{n}=\mathds{1}_{S_{n}}$ for measurable $S_{n}\subset\Gamma$, $n\geq 1$ such that $\lim_{n\to\infty}\left\lvert A\cap S_{n}\right\rvert=\int_{A}\rho(s)\,ds$ for every measurable set $A\subset\Gamma$. Then $\rho_{n}\overset{\ast}{\rightharpoonup}\rho$ in $L^{\infty}(\Gamma)$ and for fixed $k\geq 0$, $\lambda_{k}(\rho_{n})\to\lambda_{k}(\rho)$. Consequently, $\lambda_{k,\alpha}^{\bigtriangledown}=\sigma_{k,\alpha}^{\bigtriangledown}$ and $\sigma_{k,\alpha}^{\triangle}=\lambda_{k,\alpha}^{\triangle}$.
Theorem 1.3 is proven in Section 2 and has the following consequence. The inf/sup in (2)/(3) is attained by an admissible boundary partition $\Gamma=\Gamma_{S}\sqcup\Gamma_{N}$ if and only if there is a bang-bang extremal density for (7)/(8). Otherwise, there is a sequence of rapidly oscillating characteristic microstructures, $\rho_{n}$, that weak* converge to the extremal density, $\rho^{*}$, such that $\lambda_{k}(\rho_{n})\to\lambda_{k}(\rho^{*})$.
@@
$\sum_{i\in[m_{k}]}u_{i,k}^{2}\Big|_{\Gamma}(s)\begin{cases}\,\leq c,&\textrm{if }\rho_{k}^{\bigtriangledown}(s)=0,\\ \,=c,&\textrm{if }\rho_{k}^{\bigtriangledown}(s)\in(0,1),\\ \,\geq c,&\textrm{if }\rho_{k}^{\bigtriangledown}(s)=1.\end{cases}$ (10)
@@
$\sum_{i\in[m_{k}]}u_{i,k}^{2}\Big|_{\Gamma}(s)\begin{cases}\,\geq c,&\textrm{if }\rho_{k}^{\triangle}(s)=0,\\ \,=c,&\textrm{if }\rho_{k}^{\triangle}(s)\in(0,1),\\ \,\leq c,&\textrm{if }\rho_{k}^{\triangle}(s)=1.\end{cases}$ (11)
Interestingly, for a simple extremal eigenvalue (either minimizer or maximizer), the corresponding eigenfunction is constant on any interval for which the extremal density takes values between 0 and 1; see Figures 4 and 5. Theorem 1.4 is proven in Section 2. The proof uses a local variational argument relying on the Gateaux derivative of a non-simple eigenvalue; see Lemma 2.2(e). This local analysis is similar to asymptotics of Steklov eigenvalues with respect to shape perturbations [28, 29, 30, 26, 25, 27, 3].
Theorem 1.5.
(Symmetry) Consider the maximization problem (8) for $k=1$ and $\alpha\in(0,1]$. If $\Omega$ has a reflection or rotational symmetry, then there exists a maximizer $\rho_{1}^{\triangle}\in\mathcal{A}_{\alpha}$ that shares that symmetry.
Theorem 1.5 is proven in Section 2 and relies on the convexity of $\rho\mapsto\lambda_{1}^{-1}(\rho)$; see Lemma 2.2(c).
@@
Note that despite the objective function in (7) being non-convex for $k\geq 2$ (see Remark 1.2), we conjecture the minimizers in (7) to be bang-bang solutions; see Figures 1, 2, and 3.
@@
Theorem 1.8(1)–(3) is proven in Section 3. Theorem 1.8(1) is a special case of Weinstock’s inequality [31, 12]; we provide a proof for the special case where the domain $\Omega$ is fixed. Theorem 1.8(2) rules out the possibility that the constant density is the minimizer/maximizer for (7)/(8) for $k\geq 2$. Theorem 1.8(3) establishes Conjecture 1.7(1) in the asymptotic limit when $\alpha\to 1$. This is stated somewhat informally; by “in the limit as $\varepsilon\to 0$” we mean that the linearization of the problems about $\varepsilon=0$ have the stated properties. Theorem 1.8(4) shows rapidly oscillating Steklov/Neumann B.C. homogenizes and behaves like a constant density and follows directly from Theorem 1.3. Based on Conjecture 1.7(2) concerning $\rho_{k}^{\triangle}$ and Theorem 1.3, there is not a partition $\Gamma=\Gamma_{S}\sqcup\Gamma_{N}$ that attains the supremum (3) for $k\geq 1$.
In Section 4, we describe a computational approach for solving (7) and (8). We describe an efficient integral equation approach for solving the weighted Steklov eigenproblem (4) based on representing the eigenfunction in terms of a modified single layer potential. This formulation uses the weak formulation of the integral eigenvalue equation to extend the method in [1] to include boundary weights. We also describe a gradient-based optimization method for approximating the solutions to the non-convex, non-smooth extremal eigenvalue problems (7) and (8). In Section 5, these computational methods are used in several computational experiments. For the disk $\Omega=\mathbb{D}$, we observe that the extremal densities are very structured, as summarized in Conjecture 1.7.
@@
The paper is organized as follows. In Section 2, we prove Theorems 1.1, 1.3, 1.4, and 1.5. In Section 3, we further analyze the case when $\Omega$ is a disk, proving Theorem 1.8. In Section 4, we describe computational methods for weighted Steklov eigenvalues (4) as well as for approximating the solutions to the non-convex, non-smooth extremal eigenvalue problems (7) and (8). In Section 5, these computational methods are used in computational experiments for the disk. In Appendix A, we give a simple example of a Sturm-Liouville operator that has non-convex higher eigenvalues.

### v2

@@ Abstract.
Let $\Omega$ be a bounded open planar domain with smooth connected boundary, $\Gamma$, that has been partitioned into two disjoint components, $\Gamma=\Gamma_{S}\sqcup\Gamma_{N}$. We consider the Steklov-Neumann eigenproblem on $\Omega$, where a harmonic function is sought that satisfies the Steklov boundary condition on $\Gamma_{S}$ and the Neumann boundary condition on $\Gamma_{N}$. We pose the extremal eigenvalue problems (EEPs) of minimizing/maximizing the $k$-th non-trivial Steklov-Neumann eigenvalue among boundary partitions of prescribed measure. We formulate a relaxation of these EEPs in terms of weighted Steklov eigenvalues where an $L^{\infty}(\Gamma)$ density replaces the boundary partition. For these relaxed EEPs, we establish existence and prove optimality conditions. We also prove a homogenization result that allows us to use solutions to the relaxed EEPs to infer properties of solutions to the original EEPs. For a disk, we provide numerical and asymptotic evidence that the minimizing arrangement of $\Gamma_{S}\sqcup\Gamma_{N}$ for the $k$-th eigenvalue consists of $k+1$ connected components that are symmetrically arranged on the boundary. For a disk, for $k=1$, the constant density is a maximizer for the relaxed problem; we also provide numerical and asymptotic evidence that for $k\geq 2$, the maximizing density for the relaxed problem is a non-trivial function; a sequence of rapidly oscillating Steklov/Neumann boundary conditions approach the supremum value.
@@ 1. Introduction
Fix a bounded open domain $\Omega\subset\mathbb{R}^{2}$ with Lipschitz connected boundary $\partial\Omega=\Gamma$. For a disjoint partition of the boundary, $\Gamma=\Gamma_{S}\sqcup\Gamma_{N}$, we consider the Steklov-Neumann (S-N) eigenproblem:
(1a) $\Delta u$ $=0$  $\qquad\text{ in $\Omega$},$
(1b) $\partial_{n}u$ $=\sigma u$  $\qquad\text{ on $\Gamma_{S}$},$
(1c) $\partial_{n}u$ $=0$  $\qquad\text{ on $\Gamma_{N}$}.$
@@
(2) $\sigma_{k,\alpha}^{\bigtriangledown}=\inf_{\begin{subarray}{c}\Gamma_{S}\subset\Gamma\\ \left\lvert\Gamma_{S}\right\rvert=\alpha\left\lvert\Gamma\right\rvert\end{subarray}}\ \sigma_{k}(\Gamma_{S}).$
@@
(3) $\sigma_{k,\alpha}^{\triangle}=\sup_{\begin{subarray}{c}\Gamma_{S}\subset\Gamma\\ \left\lvert\Gamma_{S}\right\rvert=\alpha\left\lvert\Gamma\right\rvert\end{subarray}}\ \sigma_{k}(\Gamma_{S}).$
@@
The optimization problems (2) and (3) were posed by Nilima Nigam in the AIM Problem List (AimPL) on Steklov eigenproblems [undefx, Problem 4.1]. Steklov eigenvalues appear in many applications [undefw, undefl], so the extremal eigenvalue problems (2) and (3) are of fundamental interest. We offer a physical interpretation in terms of the analogous extremal Laplacian eigenvalue problems with mixed Dirichlet/Neumann boundary conditions (B.C.) studied by S. J. Cox and P. X. Uhlig [undefi]. These problems, for $k=1$, have the interpretation that if we are allowed to fasten a prescribed fraction of the boundary of a drum (corresponding to Dirichlet B.C.), which portion should we fasten to produce the lowest/highest bass note (first Laplacian eigenvalue)? The S-N eigenvalue problem (1) is the limit of this eigenvalue problem where the mass of the drum is concentrated at $\Gamma_{S}$ [undefy]. Thus, the optimization problem (2) for $k=1$ has the following physical interpretation: Suppose we are allowed to fasten a fixed fraction of the boundary of a drum (corresponding to Steklov B.C.), where the mass of the drum will also be concentrated. Which portion should we fasten to produce the lowest bass note?
@@
The S-N eigenvalue problem has attracted the interest of researchers in recent years. In [undefe], Bañuelos, Kulczycki, Polterovich, and Siudeja established inequalities between mixed S-N and mixed Steklov-Dirichlet eigenvalues for certain domains satisfying the weak John’s condition. These inequalities are used to obtain geometric information about the nodal sets of S-N eigenfunctions. In [undefc], Ammari, Imeri, and Nigam analyzed how small changes in the S-N partition move the spectrum and used the resulting shape-derivative formulas to design a boundary-integral method that tunes mixed S-N conditions to hit target resonances. They also treated multiple eigenvalues and validated the asymptotics numerically on canonical shapes. In [undefd], Arias-Marco, Dryden, Gordon, Hassannezhad, Ray, and Stanhope proved a sharp isoperimetric inequality for mixed Steklov eigenvalues on surfaces by relating S-N and Steklov–Dirichlet spectra, developed full asymptotics for mixed problems under mild interface regularity, and studied symmetry-constrained maximization, showing that the upper bounds converge to the Hersch-Payne-Schiffer asymptotics as $k\to\infty$. In [undefm], Grebenkov gave a “small-window” asymptotic analysis for the mixed S-N problem (framed via an integral-operator reformulation). The asymptotics are checked numerically for an arc on a disk and a spherical cap on a ball, and are linked to diffusion-controlled reactions (mean first-reaction times and reactivity-dependent corrections). In [undefg], Basak and Verma studied Steklov and mixed S-N spectra on domains with a spherical hole. They obtained an explicit expression for the second nonzero Steklov eigenvalue on an annulus and derived sharp upper bounds for the first $n$ nonzero Steklov eigenvalues on 4–fold symmetric domains in $\mathbb{R}^{n}$, with analogous bounds for the mixed problem; examples show the symmetry hypothesis is essential.
@@
(4a) $\Delta u$ $=0$  $\qquad\text{ in $\Omega$},$
(4b) $\partial_{n}u$ $=\lambda\rho u$  $\qquad\text{ on $\Gamma$}.$
@@
(5) $\lambda_{k}=\min_{E_{k}\subset H^{1}(\Omega)}\ \max_{u\in E_{k}\setminus\{0\}}\ \frac{\int_{\Omega}\left\lvert\nabla u\right\rvert^{2}dx}{\int_{\Gamma}\rho u^{2}\,ds}.$
This weighted Steklov eigenproblem is of interest in its own right, as a conformal mapping applied to the Steklov eigenvalue problem will introduce a boundary weight [undefah, undefa, undefaa].
@@
(6) $\mathcal{A}_{\alpha}\coloneqq\left\{\rho\in L^{\infty}(\Gamma)\colon 0\leq\rho(s)\leq 1\textrm{ for a.e. }s\in\Gamma,\,\int_{\Gamma}\rho(s)\,ds=\alpha\left\lvert\Gamma\right\rvert\right\}$.
@@
(7) $\lambda_{k,\alpha}^{\bigtriangledown}=\inf_{\rho\in\mathcal{A}_{\alpha}}\ \lambda_{k}(\rho),$
@@
(8) $\lambda_{k,\alpha}^{\triangle}=\sup_{\rho\in\mathcal{A}_{\alpha}}\ \lambda_{k}(\rho).$
Clearly, we have that $\lambda_{k,\alpha}^{\bigtriangledown}\leq\sigma_{k,\alpha}^{\bigtriangledown}\leq\sigma_{k,\alpha}^{\triangle}\leq\lambda_{k,\alpha}^{\triangle}$. Note that these are all finite, since by the Hersch-Payne-Schiffer inequality [undefo, undefk]
(9) $\lambda_{k}(\rho)\leq\frac{2\pi}{\alpha\left\lvert\Gamma\right\rvert}k,\qquad\forall\rho\in\mathcal{A}_{\alpha}.$
@@
(Existence) For every $k\in\mathbb{N}$ and $\alpha\in(0,1]$, there exists $\rho_{k}^{\bigtriangledown}\in\mathcal{A}_{\alpha}$ that attains the infimum in (7) and there exists $\rho_{k}^{\triangle}\in\mathcal{A}_{\alpha}$ that attains the supremum in (8).
Theorem 1.1 is proven in Section 2. The statement follows from the continuity of $\rho\mapsto\lambda_{k}(\rho)$ on $\mathcal{A}_{\alpha}$ and the compactness of $\mathcal{A}_{\alpha}$ with respect to the weak topology on $L^{\infty}(\Gamma)$.
@@
The map $\rho\mapsto\lambda_{k}^{-1}(\rho)$ is not convex for all $k\geq 1$; see [undefq, Figure 1]. This seems to be misunderstood in the literature for eigenvalues of other elliptic operators. In Appendix A, we give a simple example of a Sturm-Liouville operator that has non-convex higher eigenvalues.
@@
(Homogenization) Suppose $\rho\in L^{\infty}(\Gamma)$ and $\rho_{n}=\mathds{1}_{S_{n}}$ for measurable $S_{n}\subset\Gamma$, $n\geq 1$ such that $\lim\limits_{n\to\infty}\left\lvert A\cap S_{n}\right\rvert=\int_{A}\rho(s)\,ds$ for every measurable set $A\subset\Gamma$. Then $\rho_{n}\overset{\ast}{\rightharpoonup}\rho$ in $L^{\infty}(\Gamma)$ and for fixed $k\geq 0$, $\lambda_{k}(\rho_{n})\to\lambda_{k}(\rho)$. Consequently, $\lambda_{k,\alpha}^{\bigtriangledown}=\sigma_{k,\alpha}^{\bigtriangledown}$ and $\sigma_{k,\alpha}^{\triangle}=\lambda_{k,\alpha}^{\triangle}$.
Theorem 1.3 is proven in Section 2 and has the following consequence. The inf/sup in (2)/(3) is attained by an admissible boundary partition $\Gamma=\Gamma_{S}\sqcup\Gamma_{N}$ if and only if there is a bang-bang extremal density for (7)/(8). Otherwise, there is a sequence of rapidly oscillating characteristic microstructures, $\rho_{n}$, that weak* converge to the extremal density, $\rho^{*}$, such that $\lambda_{k}(\rho_{n})\to\lambda_{k}(\rho^{*})$. Here, we say that $\rho\in\mathcal{A}_{\alpha}$ is a “bang-bang” density if $\rho=\mathds{1}_{S}$ for a measurable set $S\subseteq\Gamma$.
@@
(10) $\sum_{i\in[m_{k}]}u_{i,k}^{2}\Big|_{\Gamma}(s)\begin{cases}\,\leq c,&\textrm{if }\rho_{k}^{\bigtriangledown}(s)=0,\\ \,=c,&\textrm{if }\rho_{k}^{\bigtriangledown}(s)\in(0,1),\\ \,\geq c,&\textrm{if }\rho_{k}^{\bigtriangledown}(s)=1.\end{cases}$
@@
(11) $\sum_{i\in[m_{k}]}u_{i,k}^{2}\Big|_{\Gamma}(s)\begin{cases}\,\geq c,&\textrm{if }\rho_{k}^{\triangle}(s)=0,\\ \,=c,&\textrm{if }\rho_{k}^{\triangle}(s)\in(0,1),\\ \,\leq c,&\textrm{if }\rho_{k}^{\triangle}(s)=1.\end{cases}$
The constant $c>0$ in the optimality conditions (10) and (11) depends on the choice of eigenbasis $\{u_{i,k}\}_{i\in[m_{k}]}$; see the proof in Section 2 for details. Interestingly, for a simple extremal eigenvalue (either minimizer or maximizer), the corresponding eigenfunction is constant on any interval for which the extremal density takes values between 0 and 1; see Figures 4 and 5. Theorem 1.4 is proven in Section 2. The proof uses a local variational argument relying on the Gateaux derivative of a non-simple eigenvalue; see Lemma 2.2(e). This local analysis is similar to asymptotics of Steklov eigenvalues with respect to shape perturbations [undefae, undefaf, undefag, undefac, undefab, undefad, undefb].
@@
Note that despite the objective function in (7) being neither convex nor concave (see Remark 1.2), we conjecture the minimizers in (7) to be bang-bang solutions; see Figures 1, 2, and 3.
@@
Theorem 1.7(2)–(3) is proven in Section 3. Theorem 1.7(1) is a special case of Weinstock’s inequality [undefah, undefk]. Theorem 1.7(2) rules out the possibility that the constant density is the minimizer/maximizer for (7)/(8) for $k\geq 2$. Theorem 1.7(3) establishes Conjecture 1.6(1) in the asymptotic limit when $\alpha\to 1$. This is stated somewhat informally; by “in the limit as $\varepsilon\to 0$” we mean that the linearization of the problems about $\varepsilon=0$ have the stated properties. Theorem 1.7(4) shows rapidly oscillating Steklov/Neumann B.C. homogenizes and behaves like a constant density and follows directly from Theorem 1.3. Based on Conjecture 1.6(2) concerning $\rho_{k}^{\triangle}$ and Theorem 1.3, there is not a partition $\Gamma=\Gamma_{S}\sqcup\Gamma_{N}$ that attains the supremum (3) for $k\geq 1$.
In Section 4, we describe a computational approach for solving (7) and (8). We describe an efficient integral equation approach for solving the weighted Steklov eigenproblem (4) based on representing the eigenfunction in terms of a modified single layer potential. This formulation uses the weak formulation of the integral eigenvalue equation to extend the method in [undef] to include boundary weights. We also describe a gradient-based optimization method for approximating the solutions to the non-convex, non-smooth extremal eigenvalue problems (7) and (8). In Section 5, these computational methods are used in several computational experiments. For the disk $\Omega=\mathbb{D}$, we observe that the extremal densities are very structured, as summarized in Conjecture 1.6.
@@
The paper is organized as follows. In Section 2, we prove Theorems 1.1, 1.3, and 1.4. In Section 3, we further analyze the case when $\Omega$ is a disk, proving Theorem 1.7. In Section 4, we describe computational methods for weighted Steklov eigenvalues (4) as well as for approximating the solutions to the non-convex, non-smooth extremal eigenvalue problems (7) and (8). In Section 5, these computational methods are used in computational experiments for the disk. In Appendix A, we give a simple example of a Sturm-Liouville operator that has non-convex higher eigenvalues.